OpenStudy (anonymous):

The cost of 3 notebooks and four pencils is $8.50. The cost of 5 notebooks and 8 pencils is $14.50. Determine the cost of one notebook and the cost of one pencil. Only in algaebraic solution. I guess it is 3n +4p =8.50 5n+ 8p=14.50 Then I am stuck

OpenStudy (zpupster):

lets try and eliminate one variable by adding the equations together if we multiply the first equation by -2 we will get -8p - 8p +8p=0p ok lets do that -2(3n +4p=8.50) -6n-8p =-17.00added to 5n+8p = 14.50 leaves us with -n= -2.50 or n=2.50 pluggung in one equation 3(2.50) +4P = 8.50 7.50 +4P = 8.50 subtract 7.50 from both sides 7.50-7.50)+ 4P = 8.50-7.50 4P = 1.00 divide both by 4 P=.25 you can do a check with the other equation

OpenStudy (anonymous):

Ok thank you soo much!